# Change point analysis of high-dimensional data using random projections

### v1

@@ 1 Introduction
Change point detection plays a crucial role in science by identifying shifts in the underlying dynamics of a system. While change point analysis has been extensively studied in low-dimensional settings (e.g., Csörgö and Horváth (1997); Aue and Horváth (2013); Horváth and Rice (2014); Aminikhanghahi and Cook (2017)), recent advances in the ability to collect and store high-dimensional data introduce new challenges. For instance, functional magnetic resonance imaging data in neuroscience (Aston and Kirch, 2012b), S&P500 companies’ stock prices in finance (Horváth et al., 2014; Jirak, 2015), and daily temperatures over decades in climate studies (Aue et al., 2018) can all be high-dimensional. Detecting the precise timing of dynamic shifts in such high-dimensional data has increasingly attracted attention in the literature. See Liu et al. (2022) for an extensive review of existing change point detection methods across different high-dimensional settings.
@@
Current change point detection methods for high-dimensional data can be broadly divided into two categories: projection-based approaches and fully functional methods. Projection-based approaches rely on classical univariate change point tools, such as CUSUM (Page, 1954, 1955) with no or minimal modifications and involve some form of dimension reduction. The first type reduces dimensions before applying a CUSUM test. The Functional Principal Component (FPC) analysis projects onto a few leading directions that capture the majority of variance (Berkes et al., 2009; Aue et al., 2009; Aston and Kirch, 2012a) while others seek a single projection direction aligned the most with the change (Wang and Samworth, 2018; Aston and Kirch, 2018). CUSUM-type statistics are then applied to the projected data. The other type applies the CUSUM tests component-wise before the dimension reduction. The CUSUM statistics are then aggregated using sum, max, or other suitable norms (Bai, 2010; Horváth and Hušková, 2012; Chan et al., 2013; Jirak, 2015; Cho, 2016; Enikeeva and Harchaoui, 2019; Liu et al., 2020; Yu and Chen, 2021; Wang et al., 2022; Wang and Feng, 2023). While these approaches leverage well-established univariate tools, reducing the data to one or a few dimensions may loose information if the projection does not fully capture the structure of the change. For instance, the FPC-based technique cannot consistently detect the change if the mean break function is orthogonal to the direction of FPCs (Aue et al., 2018).
Fully-functional methods (Horváth et al., 2014; Aue et al., 2018; Dette et al., 2020) are relatively free of such projection-induced limitations. However, they assume the underlying signal is sufficiently smooth, which may limit their applicability to certain types of data. These methods require a smoothing step since observations are often only on discrete grids, and improper smoothing can negatively affect results. For example, using too few basis functions for densely observed data may remove high-frequency variations (Jiao et al., 2022), while using too many basis functions can produce excessively large covariance matrices, leading to slow computation.
@@
In this paper, we propose a novel projection-based change point detection algorithm for high-dimensional data with a single change point, based on random projections. Random projection is a popular dimension-reduction tool (Bingham and Mannila, 2001; Dasgupta, 2013). While Aston and Kirch (2018) consider random projections in the context of change point detection for high-dimensional data, only one random projection is considered, which is proposed to be used as the lower benchmark when measuring efficiency of change point detection methods. However, random projections cannot capture the changes in the original dimension unless large enough numbers of them are involved - one random projection is certainly not enough. This paper proposes to aggregate across multiple random projections. An ensemble of many random projections and proper aggregation captures the change in the original dimension, avoiding the possible loss of information that is common in other projection-based approaches. Our method does not suffer from the limitations of the fully functional approaches, either, without requiring the smoothness assumption nor the basis functions. Our approach may not fully exploit the dimensionality reduction property of random projections, as it can require the same or even a greater number of projections than the original dimension (Lee et al., 2005). Nevertheless, our paper shall demonstrate that random projections still provide a valuable framework for change analysis.
@@
Adapting random projections to change point analysis involves many choices: projection directions and their number, the univariate change point tests, and the aggregation procedure. A guideline of such choices are provided through extensive simulations, where size and power of global tests as well as accuracy of location estimates are explored. For the univariate change point tests, cumulative sum (CUSUM) tests (Page, 1954) and its variants are considered in our simulations. Aggregation methods include the classical approaches, such as Bonferroni’s correction (Bonferroni, 1936) and Benjamini–Hochberg procedure (Benjamini and Hochberg, 1995), as well as more recently proposed $p$-value combination techniques such as harmonic mean p (Wilson, 2019). In addition, random projections may introduce high variability in the location estimation. We propose to repeat the random projection-CUSUM-aggregation procedure multiple times for stability.
@@
The advantages of the proposed method can be summarized as follows. First, using random projections, our method is easy to implement, taking advantage of the rich literature and computational ease of univariate time series. Second, compared to other projection-based methods, which heavily rely on the quality of selected directions, our method does eliminates the need for exploration of optimal projection direction and the possible information loss due to dimension reduction. Third, our method does not rely on the functional framework, which requires smooth functions.
@@ 2.1 Step 1: RP Step
As Achlioptas (2003) explains, a “threefold speedup” can be achieved using this sparse $\mathbb{d}$, which randomly drops two-thirds of the original data. To further reduce computational cost, Li et al. (2006) propose “very sparse random projections” which requires only $1/b$ of the original data to be processed, where $b\in\mathbb{Z}$ and $b\gg 3$. While a larger $b$ speeds the computation up significantly, Li et al. (2006) notice that higher sparsity increases the variance of the pairwise distances of projected data, pointing out that using $b<3$ can achieve smaller variability than a large $b$. In our approach, we follow Achlioptas (2003), setting $b=3$, to balance between computational cost and stability. For the number of projections, a theoretical bound is provided in the Johnson-Lindenstrauss (JL) lemma (Johnson and Lindenstrauss (1984)), which aims to preserve pairwise distances. However, Bingham and Mannila (2001) achieve satisfactory empirical performance using far fewer projections than the JL bound suggests. Our simulation results in Subsection 3.1 with sample size $n=50$ also reveal that using a relatively small number of projections, $k=200$, can already achieve high accuracy. The larger the signal-to-noise ratio is, the smaller the number of random projections is needed.
@@ 3 Simulation
Following Aue et al. (2018)’s setting, $D=21$ and $n=50$. The values of $SNR$ are 0, 0.1, 0.2, 0.3, 0.5, 1, and 1.5, with $SNR=0$ indicating $\delta(s)=0$ ($H_{0}$), while $SNR>0$ represent alternative cases. The break function $\delta(s)$ is constructed with $m=1$, 55, or $20$. When discretization of the functional data is needed for our RP method or for the method in Wang and Samworth (2018), we evaluate functions at equally spaced grid points $\frac{j}{p}\in[0,1]$, $j=1,\ldots,p$, with $p=101$. All results are based on 1000 simulations unless otherwise specified.
@@ 3.3 Location estimates of the RP method in practice
The location estimates based on RP methods may be unstable due to the randomness of directions. In practice, we propose to repeat RP-BH (or RP-Bonf) method multiple times to achieve stability. We compare the quality of location estimates of RP-based methods with those of existing methods in simulations: fix a single simulated dataset, and repeat each RP-BH (or RP-Bonf) method 1000 times. We consider $k=200$ random projections and the standard CUSUM. The change point location is set at $\theta=0.25$. The magnitude of the break function is scaled by $SNR=0.5$. Figure 7 summarizes the resulting empirical distributions of the 1000 estimated change point locations using violin plots, with the red dot indicating the mode across all 1000 repetitions for RP-BH and RP-Bonf. We also report the estimated change point locations only when a 200-RP set leads to a significant adjusted p-values: $\min_{1\leq r\leq 200}p_{adj(r)}<0.05$. These methods are labeled as RP_BH_sig and RP_Bonf_sig. Figure S.23 in Supplementary Material S.3 includes those results. While RP_BH_sig and RP_Bonf_sig may have smaller variability compared to those using all 1000 change location estimates, they may miss detection some changes, especially when $m=1$.
Five additional figures similar to Figure 7 are presented as Figures S.24-S.27 in Section S.3 in the Supplementary Material. These figures are the same as Figure 7, except that they are based on different datasets. In most cases, the RP methods exhibit less variability when $m=5$ and $m=20$ than $m=1$, especially in Settings 1-2. When $m=5$ and $m=20$, though repeating the RP method can yield variable estimated locations, the mode of the RP repetitions is accurate in most cases, tending to be closer to the true location than other methods. When $m=1$, the mode of the RP repetitions is similar to the locations obtained by other methods in most datasets.
@@ 5 Conclusion
This paper propose a random projection change point detection method in high-dimensional data. By projecting high-dimensional data onto multiple random directions and applying a standard univariate change point test on each projected series, the proposed method converts a high-dimensional problem into a collection of one-dimensional tasks, which is computationally easy and offers more methods to use. The family-wise error or false discovery rate is controlled by aggregating across projections using a $p$-value combination method. Our simulation results show that the RP method has better size control, higher power, and accurate location detection when the break function is not constant in the considered cases. At the same time, the change location estimate of the RP methods may have high variability. We handle it by repeating RP methods and reporting the mode over repetitions. The real-data analysis illustrates how the proposed RP method can be used in practice. In summary, the RP method provides a computationally easy and conceptually simple approach for change point analysis in high dimensions.
@@
Supplementary Material
The supplementary material presents additional simulation results. Section S.1 presents additional results for the parameter choice, relating to Section 3.1: Subsection S.1.1 presents the results of the RP method using HAC estimator, and Subsection S.1.2 presents the results of the RP method using Horváth et al. (2020)’s variance estimator. Section S.2 presents estimated change locations using weighted CUSUM approaches. Section S.3 presents figures similar to Figure 7 in Section 3.3 but with different datasets.
Appendix S.1 Additional results for Subsection 3.1
S.1.1 Result of using the HAC variance estimator
@@
S.1.2 Result of using the variance estimator in Horváth et al. (2020)
@@
Appendix S.2 Comparison of different methods using weighted CUSUM
@@
Appendix S.3 Repeating the procedure in Subsection 3.3 using additional datasets
@@
(a) Setting 1, with $m=1,5,20$
(b) Setting 2, with $m=1,5,20$
(c) Setting 3, with $m=1,5,20$
Figure S.27: Estimated change point locations detected by repeating the methods on one dataset (Dataset 6) 1000 times. For the RP methods, the mode of the estimated locations across the 1000 repetitions is marked by a red dot. The data-generating process follows (3.2) where the standard deviation $\sigma_{g}$ follows Settings 1-3. The change point location is set at $\theta=0.25$. The magnitude of the break function is scaled by $SNR=0.5$.

### v2

@@ 1 Introduction
Change point detection plays a crucial role in science by identifying shifts in the underlying dynamics of a system. While change point analysis has been extensively studied in low-dimensional settings (e.g., Csörgö and Horváth (1997); Aue and Horváth (2013); Horváth and Rice (2014); Aminikhanghahi and Cook (2017)), modern applications increasingly generate high-dimensional vectors or functional observations. Examples arise in neuroscience through functional magnetic resonance imaging (Aston and Kirch, 2012a), in finance via panels of S&P500 stock prices (Horváth et al., 2014; Jirak, 2015), and in climate studies through long term temperature trajectories (Aue et al., 2018). Detecting the precise timing of dynamic shifts in such high-dimensional or functional data has increasingly attracted attention in the literature.
@@
Current change point detection methods for high-dimensional or functional data can be broadly divided into three categories: projection-based, fully functional, and manipulations of multivariate statistics. Projection-based approaches involve some form of dimension reduction and rely on classical univariate change point tools, such as cumulative sum (CUSUM) tests (Page, 1954, 1955). For instance, functional principal component (FPC) analysis projects the high-dimensional or functional data onto a few leading directions that capture the majority of variance (Berkes et al., 2009; Aue et al., 2009; Aston and Kirch, 2012b). Wang and Samworth (2018) seeks a single projection direction that aligns the most with the change. The CUSUM-type statistics are applied to the projected data afterwards. While these approaches leverage well-established univariate tools, reducing the data to one or a few dimensions may loose information if the projection does not fully capture the structure of the change. For instance, the FPC-based technique cannot consistently detect the change if the mean break function is orthogonal to the direction of FPCs (Aue et al., 2018). Fully-functional methods (Horváth et al., 2014; Aue et al., 2018; Dette et al., 2020) are relatively free of such projection-induced limitations. However, they assume the underlying signal is sufficiently smooth, which may limit their applicability to certain types of data. These methods require a smoothing step since observations are often only on discrete grids, and improper smoothing can negatively affect results. For example, using too few basis functions for densely observed data may remove high-frequency variations (Jiao et al., 2022), while using too many basis functions can produce excessively large covariance matrices, leading to slow computation. The last category is based on multivariate statistics, often in a CUSUM form, which are then aggregated using suitable norms or other manipulations (Bai, 2010; Horváth and Hušková, 2012; Chan et al., 2013; Jirak, 2015; Cho, 2016; Enikeeva and Harchaoui, 2019; Liu et al., 2020; Yu and Chen, 2021; Wang et al., 2022; Wang and Feng, 2023). See Liu et al. (2022) for an extensive review of existing change point detection methods in high-dimensional vector settings. The method proposed in this paper can be viewed as a combination of the projection-based method and the last category. Our method also aggregates a multivariate CUSUM statistic, but unlike existing methods, this statistic is constructed from projections rather than the original high-dimensional vectors.
@@
In this paper, we employ random projections. Random projection is a popular dimension-reduction tool (Bingham and Mannila, 2001; Dasgupta, 2000). While Aston and Kirch (2018) consider a random projection in the context of change point detection for high-dimensional data, only one random projection is considered, which is proposed as the lower benchmark when measuring efficiency of change point detection methods. However, one random projection is unlikely to capture the underlying structure of the original space unless multiple projections are involved. To address this limitation, our method aggregates across multiple random projections. An ensemble of many random projections and proper aggregation captures the change in the original dimension, avoiding the possible loss of information that is common in other projection-based approaches. Our method does not suffer from the limitations of the fully functional approaches, without requiring either the smoothness assumption nor the basis functions. By relying on projections, our method avoids complications associated with other multivariate CUSUM statistics: the covariance matrix of the multivariate CUSUM statistic does not need to be estimated, nor are tools such as self-normalization or bootstrapping required, and no assumption of independence across dimensions is needed. Our approach may not fully exploit the dimensionality reduction property of random projections, as current theoretical guarantees ensure that information in the original data can be preserved only when the number of projections is equal to or greater than the original dimension (Lee et al., 2005; Ghojogh et al., 2021). Nevertheless, we will demonstrate that random projections still provide a valuable framework for change analysis.
@@
Adapting random projections to change point analysis involves many choices: projection directions and their number, the univariate change point tests, and the aggregation procedure. A guideline of such choices is provided through extensive simulations, where size and power of global tests as well as accuracy of location estimates are explored. For the univariate change point tests, CUSUM tests and its variants are considered in our simulations. Aggregation methods include the classical approaches, such as Bonferroni’s correction (Bonferroni, 1936) and Benjamini–Hochberg procedure (Benjamini and Hochberg, 1995), as well as more recently proposed $p$-value combination techniques such as harmonic mean p (Wilson, 2019) and Cauchy combination tests (Liu and Xie, 2020). In addition, random projections may introduce high variability in the location estimation. We propose to repeat the random projection-CUSUM-aggregation procedure multiple times for stability.
@@
The advantages of the proposed method can be summarized as follows. First, using random projections, our method is easy to implement, taking advantage of the rich literature and computational ease of univariate time series. Second, compared to other projection-based methods, which heavily rely on the quality of selected directions, our method eliminates the need for exploration of optimal projection direction and the possible information loss due to dimension reduction. Third, our method does not rely on the functional framework, which requires smooth functions.
@@ 2.1 Step 1: RP Step
As Achlioptas (2003) explains, a “threefold speedup” can be achieved using this sparse $\mathbb{d}$, which randomly drops approximately two-thirds of the original data. To further reduce computational cost, Li et al. (2006) propose “very sparse random projections” which requires only $1/b$ of the original data to be processed, where $b\in\mathbb{Z}$ and $b\gg 3$. While a larger $b$ speeds the computation up significantly, Li et al. (2006) notice that higher sparsity increases the variance of the pairwise distances of projected data, pointing out that using $b<3$ can achieve smaller variability than a large $b$. In our approach, we follow Achlioptas (2003), setting $b=3$, to balance between computational cost and stability. For the number of projections, a theoretical bound is provided in the Johnson-Lindenstrauss (JL) lemma (Johnson and Lindenstrauss (1984)), which aims to preserve pairwise distances. However, Bingham and Mannila (2001) achieve satisfactory empirical performance using far fewer projections than the JL bound suggests. Our simulation results in Subsection 3.1 with sample size $n=50$ also reveal that using a relatively small number of projections, $k=200$, can already achieve high accuracy. The larger the signal-to-noise ratio is, the smaller the number of random projections is needed.
@@ 2.2 Step 2: CUSUM and Combination Step
The validity of this location estimate is anchored through the RP’s ability to capture the pairwise distance (Lee et al., 2005), even with the $L_{\infty}$ norm, as long as the number $k$ of projections is large. Recall that under the null, the random behavior of a projected data $\mathbb{y}_{r}$ using a random projection is similar to that of a component of the high-dimensional data (Theorem 3.1 of Aston and Kirch (2018)). When $\tilde{r}$ is chosen with $p^{\textit{Bonf}}_{adj(r)}$, Yu and Chen (2021) provide the rate of convergence of $\hat{z}^{*}$ to the true change location $z^{*}$ when weighted CUSUM $T_{z,r}^{w}$ (Theorem 3.4) or standard CUSUM $T_{z,r}^{s}$ (Theorem 3.5) is used.
@@ 3 Simulation
Following Aue et al. (2018)’s setting, $D=21$ and $n=50$. The values of $SNR$ are 0, 0.1, 0.2, 0.3, 0.5, 1, and 1.5, with $SNR=0$ indicating $\delta(s)=0$ ($H_{0}$), while $SNR>0$ represent alternative cases. The break function $\delta(s)$ is constructed with $m=1$, 55, or $20$. When discretization of the functional data is needed for our RP methods or for the method in Wang and Samworth (2018), we evaluate functions at equally spaced grid points $\frac{j}{p}\in[0,1]$, $j=1,\ldots,p$, with $p=101$. All results are based on 1000 simulations unless otherwise specified.
@@ 3.3 Location estimates of the RP method in practice
The location estimates based on RP methods may be unstable due to the randomness of directions. In practice, we propose to repeat RP-BH (or RP-Bonf) method multiple times to achieve stability. We compare the quality of location estimates of RP-based methods with those of existing methods in simulations: fix a single simulated dataset, and repeat each RP-BH (or RP-Bonf) method 1000 times. We consider $k=200$ random projections and the standard CUSUM. The change point location is set at $\theta=0.25$. The magnitude of the break function is scaled by $SNR=0.5$. Figure 7 summarizes the resulting empirical distributions of the 1000 estimated change point locations using violin plots, with the red dot indicating the mode across all 1000 repetitions for RP-BH and RP-Bonf. We also report the estimated change point locations only when a 200-RP set leads to a significant adjusted $p$-values: $\min_{1\leq r\leq 200}p_{adj(r)}<0.05$. These methods are labeled as RP_BH_sig and RP_Bonf_sig. Figure S.23 in Supplementary Material S.3 includes those results. While RP_BH_sig and RP_Bonf_sig may have smaller variability compared to those using all 1000 change location estimates, they may miss detecting some changes, especially when $m=1$.
Four additional figures similar to Figure 7 are presented as Figures S.24-S.27 in Section S.3 in the Supplementary Material. These figures are the same as Figure 7, except that they are based on different datasets. In most cases, the RP methods exhibit less variability when $m=5$ and $m=20$ than $m=1$, especially in Settings 1-2. When $m=5$ and $m=20$, though repeating the RP method can yield variable estimated locations, the mode of the RP repetitions is accurate in most cases, tending to be closer to the true location than other methods. When $m=1$, the mode of the RP repetitions is similar to the locations obtained by other methods in most datasets.
@@ 5 Conclusion
This paper proposes a random projection change point detection method in high-dimensional data. By projecting high-dimensional data onto multiple random directions and applying a standard univariate change point test on each projected series, the proposed method converts a high-dimensional problem into a collection of one-dimensional tasks, which is computationally easy and offers more methods to use. The family-wise error or false discovery rate is controlled by aggregating across projections using a $p$-value combination method. Our simulation results show that the RP method has better size control, higher power, and accurate location detection when the break function is not constant in the considered cases. At the same time, the change location estimate of the RP methods may have high variability. We handle it by repeating RP methods and reporting the mode over repetitions. The real-data analysis illustrates how the proposed RP method can be used in practice. In summary, the RP method provides a computationally easy and conceptually simple approach for change point analysis in high dimensions.
@@
Acknowledgement
Rho was partially supported by NIH-R15GM135806.
Supplementary Material on “Change point analysis of high-dimensional data using random projections”
The supplementary material presents additional simulation results. Section S.1 presents additional results for the parameter choice, relating to Section 3.1: Subsection S.1.1 presents the results of the RP method using HAC estimator, and Subsection S.1.2 presents the results of the RP method using Horváth et al. (2020)’s variance estimator $\hat{\sigma}^{2}_{z}=\frac{1}{n}[(\sum_{t=1}^{z}(y_{t}-\bar{y}_{z})^{2}+\sum_{t=z+1}^{n}(y_{t}-\bar{y}_{n-z})^{2})]$. Section S.2 presents estimated change locations using weighted CUSUM approaches. Section S.3 presents figures similar to Figure 7 in Section 3.3 but with different datasets.
S.1 Additional results for Subsection 3.1
S.1.1 Results of using the HAC variance estimator
@@ Acknowledgement
S.1.2 Results using the variance estimator in Horváth et al. (2020)
@@
S.2 Comparison of different methods using weighted CUSUM
@@
S.3 Repeating the procedure in Subsection 3.3 using additional datasets
@@
(a) Setting 1, with $m=1,5,20$
(b) Setting 2, with $m=1,5,20$
(c) Setting 3, with $m=1,5,20$
Figure S.23: Estimated change point locations detected by repeating the methods on one dataset (Dataset 1) 1000 times. For the RP methods, the mode of the estimated locations across the 1000 repetitions is marked by a red dot. The data-generating process follows (3.2) where the standard deviation $\sigma_{g}$ follows Settings 1-3. The change point location is set at $\theta=0.25$. The magnitude of the break function is scaled by $SNR=0.5$.